OpenStudy (anonymous):

Please help differentiate: 2 x^2 + 2 xy - 1 y^3 = 47

OpenStudy (anonymous):

4x+ x(dy/dx)- 3(dy/dx)=0 is y supposed to be a function of x, aka f(x) or is it just a variable?

OpenStudy (anonymous):

a function of x so we are trying to find dy/dx

OpenStudy (anonymous):

4x+ x(dy/dx)- 3(dy/dx)=0 solve for dy/dx 4x = 3(dy/dx)-x(dy/dx) 4x = (dy/dx)(3-x) 4x/(3-x)=dy/dx

OpenStudy (anonymous):

and how would you get dy/dx using implicit differentiation?

OpenStudy (anonymous):

do you have an x value for which you want to find he point of tangency?

OpenStudy (anonymous):

I do i just got it actually, thank you!